# MorphMark: Flexible Adaptive Watermarking for Large Language Models

### v1

@@ Abstract
Zongqi Wang1, Tianle Gu1 Baoyuan Wu2∗, Yujiu Yang1thanks: Yujiu Yang and Baoyuan Wu are co-corresponding authors. 1Tsinghua University  2The Chinese University of Hong Kong, Shenzhen 1{zq-wang24@mails, gtl23@mails, yang.yujiu@sz}.tsinghua.edu.cn 2wubaoyuan@cuhk.edu.cn
@@ 4.6.1 Different Sampling Parameters
These results indicate that MorphMarkexp performs still reliably across different sampling settings, maintaining high detection effectiveness while adapting to different decoding parameters. Results of MorphMarklinear and MorphMarklog are present in Tab. 3 and Tab. 4 of App. C.4
@@ C.5 Statistical Distribution of $P_{G}$ in C4 Dataset
C.6 Evaluation of Watermark Effectiveness in Low-Entropy Scenario
Experiment Setting:
We perform experiments on the HumanEval Chen et al. (2021). In this study, we employ MorphMark and evaluate it using the StarCoder2-3B Lozhkov et al. (2024), a widely used code generation model. Since MorphMark and EWD operate at different stages, we also investigate a combination approach where MorphMark is used for watermark generation and EWD for watermark detection.
Results:
The results of the experiments are summarized in Table 5. We use pass@1 as evaluation metric.
Method TPR@1$\uparrow$ Best F1$\uparrow$ pass@1$\uparrow$
No Watermark - - 0.2500
SWEET 0.8110 0.8926 0.1890
EWD 0.8537 0.9180 0.2073
MorphMark 0.8780 0.9320 0.2378
MorphMark + EWD 0.9207 0.9557 0.2378
The results indicate that MorphMark consistently outperforms EWD in low-entropy scenarios. Notably, the combination of MorphMark and EWD achieves the best overall performance, demonstrating that these methods can mutually enhance each other. This finding motivates future research to design adaptive generation and detection methods simultaneously.
C.7 Comparative Analysis of Z-Scores with Low-Entropy Watermarking Methods

### v2

@@ Abstract
Zongqi Wang, Tianle Gu, Baoyuan Wuthanks: Yujiu Yang and Baoyuan Wu are co-corresponding authors., Yujiu Yang11footnotemark: 1 1Tsinghua University  2The Chinese University of Hong Kong, Shenzhen 1zq-wang24@mails.tsinghua.edu.cn, gtl23@mails.tsinghua.edu.cn 2wubaoyuan@cuhk.edu.cn, 1yang.yujiu@sz.tsinghua.edu.cn
@@ 4.6.1 Different Sampling Parameters
These results indicate that MorphMarkexp performs still reliably across different sampling settings, maintaining high detection effectiveness while adapting to different decoding parameters. Results of MorphMarklinear and MorphMarklog are present in Tab. 4 and Tab. 5 of App. C.4.
@@ 4.6.2 In-Depth Analysis of $P_{G}$ Distribution
4.7 Evaluation in Low-Entropy Scenarios
To validate the performance of MorphMark in low-entropy scenarios, we conduct experiments on HumanEval Chen et al. (2021). We employ StarCoder2-3B Lozhkov et al. (2024), a widely used code generation model, as the test model. In this evaluation, we compare three methods: SWEET, EWD and MorphMarkexp. Given that MorphMark and EWD operate at different stages, we also explore a combined approach, where MorphMark is used for watermark generation, while EWD is employed for watermark detection. We use pass@1 as the evaluation metric.
Method TPR@1$\uparrow$ Best F1$\uparrow$ pass@1$\uparrow$
No Watermark - - 0.2500
SWEET 0.8110 0.8926 0.1890
EWD 0.8537 0.9180 0.2073
MorphMark 0.8780 0.9320 0.2378
MorphMark + EWD 0.9207 0.9557 0.2378
The results of the experiments are summarized in Table 3. The results indicate that MorphMark consistently outperforms SWEET and EWD in low-entropy scenarios. Notably, the combination of MorphMark and EWD achieves the best overall performance, demonstrating that these methods can mutually enhance each other. This finding motivates future research to design adaptive generation and detection methods simultaneously. In addition, we conduct an analytical experiment targeting low-entropy scenarios to investigate the mechanisms of SWEET, EWD, and MorphMark, as detailed in App C.6.
@@ C.5 Statistical Distribution of $P_{G}$ in C4 Dataset
C.6 Comparative Analysis of Z-Scores with Low-Entropy Watermarking Methods